Tier1 Operations Report 2018-02-27

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 28th February 2018

Review of Issues during the week 21st February 2018 to 28th February 2018
  • There is an outage on Castor announced for Thursday (1st March) while the database systems have Oracle patches applied.
  • The IPv6 connections, both to the RAL Core network and the 'bypass' route were successfully moved to share the IPv4 connections.This was done ahead of a change to enable IPv6 (dual stack) on the Echo Gateways which was enabled yesterday (27th).
  • AliceDisk is getting very full. Alice have been in informed and we continue to monitor.
Current operational status and issues
  • The problems with data flows through the site firewall being restricted is still present. New firewall kit has be purchased and delivered, awaiting deployment.
  • Echo Gateway Dual Stack upgrade was implemented but appears to have broken RAL FTS. Currently CERN FTS is being used as a stop-gap solution. This issues is on-going at time of writing.
Resolved Castor Disk Server Issues
  • gdss776 (d1t0) - Back in production (RO)
  • gdss762 (d1t0) - Back in production (RO)
Ongoing Castor Disk Server Issues
  • None
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
Downtime ID Hosts Start time End time Severity Description
24947 srm-biomed.gridpp.rl.ac.uk (SRM),srm-dteam.gridpp.rl.ac.uk (SRM), srm-cms.gridpp.rl.ac.uk (SRM),srm-atlas.gridpp.rl.ac.uk (SRM),srm-lhcb.gridpp.rl.ac.uk (SRM), srm-minos.gridpp.rl.ac.uk (SRM), srm-alice.gridpp.rl.ac.uk (SRM), srm-ilc.gridpp.rl.ac.uk (SRM), srm-mice.gridpp.rl.ac.uk (SRM), srm-t2k.gridpp.rl.ac.uk (SRM),srm-preprod.gridpp.rl.ac.uk (SRM), srm-cert.gridpp.rl.ac.uk (SRM), srm-na62.gridpp.rl.ac.uk (SRM), srm-snoplus.gridpp.rl.ac.uk (SRM), srm-lhcb.gridpp.rl.ac.uk (SRM.nearline), srm-cms-disk.gridpp.rl.ac.uk (SRM), srm-pheno.gridpp.rl.ac.uk (SRM), srm-solid.gridpp.rl.ac.uk (SRM) 01/03/2018 11:00 01/03/2018 17:00 SCHEDULED OUTAGE Patching Castor Neptune and Pluto Primary Environments (GI+DB PSU + OJVM PSU OCT 2017)
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.

Listing by category:

  • Echo
    • Applying minor CEPH update to fix the "backfill" bug. Increase number of placement groups and add more capacity.
  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Networking
    • Extend the number of services on the production network with IPv6 dual stack. (Done for Perfsonar, FTS3, all squids and the CVMFS Stratum-1 servers).
    • Replacement (upgrade) of RAL firewall.
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • Testing of power distribution boards in the R89 machine room is being scheduled for some time late July / Early August. The effect of this on our services is being discussed.
Open GGUS Tickets (Snapshot during morning of meeting)
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
117683 none on hold less urgent 18/11/2015 03/01/2018 Information System CASTOR at RAL not publishing GLUE 2 EGI
124876 ops on hold less urgent 07/11/2016 13/11/2017 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
127597 cms on hold urgent 07/04/2017 29/01/2018 File Transfer Check networking and xrootd RAL-CERN performance EGI
132589 lhcb in progress very urgent 21/12/2017 23/02/2018 Local Batch System Killed pilots at RAL WLCG
133619 cms in progress top priority 21/02/2018 27/02/2018 CMS_Central Workflows T1_UK_RAL Unmerged files missing WLCG
133717 cms in progress very urgent 27/02/2018 27/02/2018 CMS_Data Transfers RAL FTS3 Service: Significant Drop in Transfer Efficiency WLCG
133719 atlas in progress urgent 27/02/2018 27/02/2018 File Transfer Transfers to RAL-LCG2-ECHO failing WLCG
GGUS Tickets Closed Last week
  • None
Availability Report
2018-02-22 100 100 100 100 100 100
2018-02-23 100 100 100 100 100 100
2018-02-24 100 100 100 100 100 100
2018-02-25 100 100 100 100 100 100
2018-02-26 100 100 100 100 100 100
2018-02-27 100 100 100 100 100 100
2018-02-28 100 100 100 100 100 100
Hammercloud Test Report

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Day Atlas HC CMS HC Comment
2018/02/21 95 99
2018/02/22 98 99
2018/02/23 97 99
2018/02/24 100 99
2018/02/25 96 99
2018/02/26 100 100
2018/02/27 95 100
Notes from Meeting.